Find the missing coordinate. (-3,_____) is a solution of y = 4-3x.
Answer by algebrahouse.com(1659)

(-3, ?) y = 4 - 3x

Substitute -3 in for x, to find y.

y = 4 - 3(-3) {substituted -3 in for x}
y = 4 + 9 {multiplied -3 by -3}
y = 13 {added 4 and 9}
www.algebrahouse.com